Please help show me how to solve this quadratic equation by completing the square.
m^2 = 8m + 3

Found 2 solutions by scott8148, mathbird:
Answer by scott8148(6628)

subtracting 8m __ m^2-8m=3

adding half of the coefficient of the 1st order term, squared __ m^2-8m+(-8/2)^2=3+(-8/2)^2

m^2-8m+16=3+16 __ taking square root __ m-4=ħsqrt(19)

adding 4 __ m=4ħsqrt(19)

Answer by mathbird(71)

m^2 = 8m + 3
m^2-8m-3=0
(m)^2-2(m)(4)+(4)^2-(4)^2-3=0
(m-4)^2-16-3=0
(m-4)^2-19=0
(m-4)^2=19
ħ√(m-4)^2=ħ√19
m-4=ħ4.36
m=ħ4.36+4
m=8.36&-0.36